OpenStudy (anonymous):

3. Riley finds out that Cooper’s Storage and Shipping Company is working with a local business to package some office supplies. Some of the supplies are packed inside a cube-shaped box with side lengths of 4.5 in. These boxes are then packed into a shipping box with dimensions of 18 in. × 9 in. × 4.5 in. How many boxes of office supplies can be packed into the larger box for shipping? Show your work.

OpenStudy (danjs):

First need to find the volume of the small box and the large shipping box

OpenStudy (danjs):

Volume of Cube = (side length)^3 Volume of Box = length x width x height

OpenStudy (anonymous):

729

OpenStudy (danjs):

Then the number of cubes that will fit in the box is \[\frac{ Volume ~of~box }{ ~volume ~of ~cube }\]

OpenStudy (danjs):

what is the volume of the small box

OpenStudy (danjs):

volume of cube = (side length)^3 = side length x side length x side length

OpenStudy (anonymous):

91

OpenStudy (danjs):

divide the larger volume 729 by the smaller volume 91.125

OpenStudy (anonymous):

8.0098

OpenStudy (danjs):

729/91.125 = 8

OpenStudy (danjs):

so 8 boxes will fit in the larger box

OpenStudy (anonymous):

you're a good teacher

OpenStudy (danjs):

thanks :)